### v2

# A Study of Commonsense Reasoning over Visual Object Properties

Abstract
Inspired by human categorization, object property reasoning involves identifying and recognizing low-level details and higher-level abstractions. While current visual question answering (VQA) studies consider multiple object properties, such as size, they typically blend perception and reasoning and lack representativeness in terms of reasoning and image categories, making it unclear whether and how vision-language models (VLMs) abstract and reason over depicted objects. To this end, we introduce a systematic evaluation framework comprising images of three representative types, three reasoning levels of increasing complexity, and four object property dimensions, informed by prior work on common sense. We develop a procedure to instantiate this framework in two VQA object reasoning benchmarks: OPTICS-CNT, comprising 360 images paired with 1,080 multi-level, count-based questions, and OPTICS-CMP, with 2.1k comparison questions. Experiments with 12 state-of-the-art VLMs in zero-shot settings reveal significant limitations relative to humans, with the best-performing model achieving below 40% counting and 70% comparison accuracy. VLMs struggle particularly with photographic images, counterfactual reasoning, physical and functional properties, and higher counts. We make the OPTICS benchmark data and code available to support future work on scalable benchmarking methods, generalized annotation guidelines, and advanced reasoning VLMs.
1 Introduction
Object representation and reasoning is one of the core systems of human cognition (Spelke and Kinzler, 2007). By abstracting object properties and interactions, humans, including infants with limited real-world knowledge and experience, represent their environment and perform inference (Lipton and Spelke, 2004). Object representation and reasoning in humans are driven by spatio-temporal principles of cohesion (objects move as connected, bounded wholes), continuity (objects move along connected, unobstructed paths), and contact (objects do not interact at a distance). Humans perceive object boundaries, represent object shapes that move (partially) out of view, and predict object movement (Aguiar and Baillargeon, 1999; Leslie and Keeble, 1987). Importantly, object perception and inference is not merely descriptive; it involves a multi-stage process from low-level recognition, through categorization, to identification and reasoning Fields (2016). For instance, to count how many mammals are visible in Figure 1-top, humans recognize the image objects, decide which of them fit the constraint (type) of a mammal, and count those.
Figure 1: VQA in OPTICS-CNT, with questions about different object properties of varying reasoning complexity on three types of images from different domains.
Category VQAv2 CLEVR TallyQA SCIENCEQA GQA C-VQA CFMM OPTICS
Question Counting
Comparison
Physical
Object Taxonomic
Property Functional
Relational
Image Type Real
Animated
AI-Generated
Reasoning Direct Recognition
Property Inference
Counterfactual
Table 1: Comparison of OPTICS to closely related existing benchmarks: VQAv2 Goyal et al. (2017), CLEVR Johnson et al. (2017), TallyQA Acharya et al. (2018), SCIENCEQA Lu et al. (2022), GQA Hudson and Manning (2019), C-VQA Zhang et al. (2024b), and CFMM Li et al. (2024).
Recent work on the connection between perception and reasoning has used visual question answering (VQA) in natural language as a diagnostic task for vision-language models (VLMs). Existing VQA benchmarks evaluate recognition of basic visual elements such as physical attributes or taxonomic category membership Yuan et al. (2021); Tong et al. (2024), inference of relational properties Lu et al. (2022); Hudson and Manning (2019), and counterfactual reasoning Frohberg and Binder (2021); Yu et al. (2023); Li et al. (2024). As objects are central components of images, solving these benchmarks relies on extracting and reasoning over physical and abstract object properties  Johnson et al. (2017); Hudson and Manning (2019); Acharya et al. (2018).
Current VQA evaluation exhibits two key limitations. The first is the lack of a delineation between subtasks for object perception and reasoning, i.e., object detection, inference, and counterfactual reasoning. While some studies have identified small object sizes and peripheral locations as causes of failure in VLMs (Zhang et al., 2025, 2024a), the lack of subtask separation hinders comprehensive diagnosis of their strengths and weaknesses. Second, we note a limited representativeness of the covered evaluation dimensions, usually focusing on basic features such as shape and color (Goyal et al., 2017; Antol et al., 2015).  Jiang et al. (2024)’s recent systematic evaluation of abstract visual reasoning covers different types, configurations, and shapes. Given its focus on manipulating simple shapes in synthetic environments with minimal noise, the ability of VLMs to perceive and reason over real-world objects across image types remains underexplored.
We address this gap in the understanding of the capabilities of current models with a comprehensive, fine-grained evaluation of the ability of VLMs to perceive and reason given a representative range of object properties, reasoning complexities, and image types. We make three contributions (Figure 1): 1. A systematic evaluation frameworkfor object property reasoning, consisting of three image types, three levels of reasoning complexity, and four dimensions of commonsense knowledge. 2. A semi-automatic procedurefor dataset creation, resulting in two VQA benchmarks based on our framework: OPTICS-CNT with 1.1k high-quality counting questions about 360 images, and OPTICS-CMP with 2.1k comparison questions about image pairs. [^1] OPTICS stands for Object Property Reasoning Tasks for Evaluating Image-based Common Sense. 3. An extensive experimental analysisof state-of-the-art (SotA) VLMs on OPTICS, revealing fine-grained limitations of current models for photographic images, counterfactual reasoning, functional knowledge, and higher counts.
2 Related Work
Object Properties in VQA. VQA has long served as a diagnostic task for evaluating the perceptual and reasoning capabilities of AI systems. Early efforts Antol et al. (2015); Goyal et al. (2017) emphasized general visual recognition of objects, colors, and basic attributes, while subsequent benchmarks Singh et al. (2019); Mathew et al. (2021); Lu et al. (2022) targeted domain-specific context. Although many benchmarks across domains Johnson et al. (2017); Hudson and Manning (2019); Acharya et al. (2018) require reasoning over object properties, none of them provides a systematic separation of object property dimensions. Compared to previous VQA benchmarks, our framework covers four distinct object dimensions based on commonsense resources Speer et al. (2017); Ilievski et al. (2021); Kurtz and Silliman (2021): physical, taxonomic, functional, and relational, and enables fine-grained assessment of object-level reasoning as visualized in Table 1. Moreover, while both counting and comparison questions have been relatively well-covered in prior work, our evaluation enables compositional evaluation by comparing counts across images rather than object properties (e.g., which object is larger Li et al. (2024)).
Visual Reasoning Complexity. Human cognition is inherently compositional and integrates multiple scene aspects into higher-level reasoning, with complexity progressing from direct recognition to property-level inference, and counterfactual reasoning Hoffman and Richards (1987); Xu et al. (2021). Initial direct recognition is limited to recognition of basic visual elements such as physical attributes or taxonomic category membership Yuan et al. (2021); Tong et al. (2024). Inference builds on recognition, targets higher-level functional or relational properties, and requires multi-step abstraction beyond surface-level features Lu et al. (2022); Hudson and Manning (2019). Counterfactual reasoning involves reasoning about hypothetical changes made in inputs. This type of reasoning is most challenging, as it requires understanding and adapting to hypothetical, altered, or out-of-context scenarios Frohberg and Binder (2021); Yu et al. (2023); Li et al. (2024). Benchmarks that target different types of reasoning exist (Table 1), yet none systematically test reasoning levels simultaneously. Inspired by compositional evaluation frameworks in abstract visual reasoning Jiang et al. (2024), OPTICS is the first to encompass all three reasoning types and go beyond previous work to study VLM performance across three image types, thereby enabling a more comprehensive assessment of models’ perception and reasoning abilities. To support fair evaluation, we focus on counting and comparison questions in line with prior benchmarks Acharya et al. (2018); Johnson et al. (2017); Li et al. (2024).
3 A Commonsense Framework for Evaluating Object Property Reasoning
We devise a systematic framework for rigorous evaluation of VLMs, integrating object property dimensions (physical, taxonomic, functional, relational), levels of reasoning complexity (recognition, property inference, counterfactual), and diverse image types (photographic, animated, AI-generated). We next detail its three components.
3.1 Object Property Dimensions
Drawing on prior research Ilievski et al. (2021); Tandon et al. (2017) on representing commonsense knowledge about properties and their relations, our framework distinguishes four object property dimensions relevant for counting and comparison VQA. Each dimension captures a distinct facet of human conceptualization and reasoning about object attributes and relations in visual scenes.
Physical knowledge refers to the dynamics of physical systems based on observable phenomena and fundamental principles (McCloskey et al., 1983) (e.g., circular objects in Figure 1). Physical properties, such as object qualities, materials, and part-whole relations, are prevalent in commonsense knowledge representation (Tandon et al., 2017; Fleming, 2017). Our framework enhances the list of physical properties in prior VQA studies, such as shape, color, and size (Table 1), with more complex attributes like the material (wood, metal), state of the object (liquid, solid), and structural characteristics of the object – including part-whole relationships such as has wheels.
Taxonomic knowledge captures an object’s semantic category or class membership, typically expressed as is-a relations (Ilievski et al., 2021). Examples in our framework include broad ontological groupings, e.g., biological (e.g., mammals, reptiles), artifact (e.g., furniture, tools), and food (e.g., fruits, vegetables) categories. Unlike physical attributes, taxonomic properties may not be visible from visual input alone: to determine how many mammals are present in an image (Figure 1), visual identification of species must be combined with knowledge of their taxonomic classification. The taxonomic dimension aligns with taxonomic semantic systems from cognitive science and domain-specific knowledge representation, which group entities based on shared features and categorical similarity (Mirman et al., 2017).
Functional properties express attributes such as utilities and capabilities of an object (e.g., means of transportation in Figure 1), affordances as actions an object receives (e.g., breakable, foldable), and operational dependencies for an object to function (e.g., electricity, battery). This dimension primarily aligns with the utility and causal relations identified in commonsense knowledge frameworks (Speer et al., 2017; Heindorf et al., 2020; Ilievski et al., 2021). Questions about the utility of objects have been included in VQA studies (Table 1): for example, the OK-VQA benchmark (Marino et al., 2019) tests for knowledge about object utility, e.g., a broom is for sweeping.
Relational knowledge captures how objects interact, how they can be contextually grouped, and how they are situated relative to each other within a visual context. Attributes include spatial relationships (as in "How many objects can be seen hanging from the wall?") and contextual grouping relations (as in "How many couples are visible?"), inspired by commonsense knowledge organization (Ilievski et al., 2021). Its emphasis on relations between multiple object instances complements the focus on properties of single objects in taxonomic and part-whole physical properties.
3.2 Reasoning Complexity
The four object property dimensions can support questions with varying reasoning complexity. Our three levels of reasoning complexity: direct recognition, property inference, and counterfactual reasoning are exemplified in Figure 1. We expect that the progression from direct recognition to counterfactual reasoning requires an increasingly high level of abstraction to answer a question, and a shifting emphasis from perception to reasoning, thus increasing in difficulty (Xu et al., 2021).
Direct Recognition questions ask about features that can be detected by observation and general categorical knowledge. Questions at this level can be answered through an atomic step of perception and reside at the physical and taxonomic property dimensions. For example, the top question in Figure 1 only requires identifying land animals and categorizing them as mammals.
Property Inference questions target the functional and relational property dimensions, which require deeper abstraction, introducing multi-step reasoning. It studies the ability of models to generalize beyond surface-level features by focusing on codependencies among objects and underlying relations across different image regions. For example, in Figure 1, the middle question requires identifying a bicycle, a car, and a plane, categorizing them as vehicles, and then reasoning about their utility in transportation.
Counterfactual Reasoning questions probe the ability to reason about hypothetical, altered, or out-of-context scenarios. Counterfactual questions span any of the four property dimensions. In addition to visual recognition and inference, counterfactual questions also require commonsense knowledge and flexible abstraction (e.g., creating a contextualized procedure that identifies remaining circular objects besides half of the clocks in the bottom question of Figure 1), thus forming the most complex reasoning level in our framework.
3.3 Image Types
To study VLM generalization across visual domains, we cover various image types, expecting that accuracy varies (often worsens) when moving from synthetic to real-world images (Johnson et al., 2017; Hudson and Manning, 2019).
Photographic images represent naturalistic settings, featuring complex textures, lighting variations, and cluttered scenes. They are realistic, i.e., all depicted objects are found in the real world. Photographs present challenges such as occlusions, viewpoint variation, and natural ambiguity, requiring an effective visual grounding.
Animated images are simplified or stylized and contain minimal visual noise, as well as limited levels of detail when compared to photographic images. Humans typically create animated images in digital form, and their objects are not necessarily realistic. Object boundaries in animated images may be easier to detect, but exaggerated representations may pose a novel challenge.
AI-generated images include both photorealistic and stylized images, created by prompting generative models. They test robustness to domain shifts by deviating from visual conventions such as real-world physics and object proportions. AI-generated images typically feature simpler compositions and include implausible objects, such as half-animals or floating glasses.
Figure 2: OPTICS-CNT and OPTICS-CMP construction pipeline: image collection, candidate QA formulation, quality assurance, and pairwise combination. n and m are number of images and QAs, respectively.
4 Benchmark Construction
Our dataset curation follows a four-phase pipeline (Figure 2). We collect images from public sources corresponding to the three image types. Then, we ask an annotator to formulate candidate question-answer pairs for all images, inspired by AI-generated questions. Then, the questions and answers undergo additional rounds of human refinement for quality assurance, which results in OPTICS-CNT. OPTICS-CMP is then systematically constructed from samples in OPTICS-CNT. Finally, we present statistics about the two benchmarks. We detail each step below.
Image Collection. To ensure diagnostic quality and mitigate data contamination risks from existing benchmarks (Dodge et al., 2021), we collect previously unsourced images. We define 26 domains grouped into three categories based on the MacGyver dataset (Tian et al., 2025) (see Table 6 for details). We use these domains as queries across all three image types, enabling fair comparisons of images from comparable scenes. We collect public images by querying a variety of sources: Google Images, Unsplash, and Freepik for photographic and animated images; GPT-4o and Grok3 for AI-generated ones (see Table 5 for example generative AI prompts). We explicitly exclude diagrammatic images (Lu et al., 2022). In total, we collect 360 images (120 per type).
Formulation of Candidate QA-pairs. We formulate initial question-answer pairs using a semi-automatic procedure. Namely, we prompt a multimodal large-language model (MLLM) to generate three counting questions (one per reasoning level) following the framework defined in the previous section. The MLLM is selected at random from GPT-4o (OpenAI et al., 2024), Claude 3.7 Sonnet (Anthropic, 2025), and Gemini 2.0 Flash (Google et al., 2025a). The resulting 1,080 question-answer pairs are then used by five human annotators, who are instructed to revise or replace them to ensure high quality. Namely, humans ensure that the question is precise and fits the restrictions of our evaluation framework, and that the answer count is not too high (we cap the answers at 10). Each of the annotators is responsible for 72 images. When possible, the annotators are instructed to abstract over the specific object types, e.g., prioritizing “how many objects are made of rubber?” over “how many tires are made of rubber?”. In practice, the annotators rewrote a vast majority of the questions and nearly all answers, as many MLLM-generated questions were ambiguous or repetitive, and most answers were incorrect.
Quality Assurance and Consolidation of OPTICS-CNT. Human annotators may introduce controversial assumptions and biases. To ensure a broader consensus and catch ambiguous questions, we perform two rounds of quality assurance. First, each of the 360 images is randomly assigned to an annotator (from the same pool) who did not annotate the QA pair in the previous phase. This annotator is provided with the question and the image, and asked to provide an answer. Comparing the two annotators’ answers for each question, we observed that they agreed on 756 answers (70%). We identified two key disagreement sources: 1) Semantic ambiguity: Uncertainty in the definition, including objects that partially satisfy a property (e.g., a knife that is partly metal) or objects with unclear definitions (e.g., a container). 2) Visual ambiguity: Uncertainty in the visual information, such as deciding whether to count a partially visible chair. We addressed these disagreements by another round of quality assurance, where a third, different annotator looked at the question, answer, and comments and attempted to resolve the disagreement by rephrasing the questions to make them more precise, or if that was challenging, to replace them with a new QA pair from the same reasoning complexity level. This results in OPTICS-CNT.
Construction of OPTICS-CMP. To introduce an evaluation axis beyond exact counting and diversify question types, we devise OPTICS-CMP, which assesses the ability to make relative judgments between images. OPTICS-CMP is derived from annotated VQA samples in OPTICS-CNT. We first group questions by object properties and reasoning complexity, then pair them using semantic similarity Reimers and Gurevych (2019). For each pair, the corresponding images are resized and concatenated. Then, a comparison question is generated using the template in Figure˜3, and an answer is automatically derived from the OPTICS-CNT labels (see example in Figure˜9).
Let AA be the answer to left question for the left image. Let $B$ be the answer to right question for the right image. Which is greater, AA or $B$?
Figure 3: OPTICS-CMP question template.
Figure 4: OPTICS-CNT question composition.
Benchmark Composition and Analysis. OPTICS-CNT integrates the images, questions, and final ground truth answers in a dataset containing 1,080 questions about 360 high-quality images of photographic, animated, and AI-generated types in equal proportions. The overall distribution across object properties is also relatively balanced (Figure 4). We observe a larger portion of taxonomic over physical questions in the direct recognition category, and similarly, slightly more relational than functional questions in the inference category. Curiously, when creating counterfactual questions, annotators prioritized physical and taxonomic over relational and functional knowledge. Meanwhile, OPTICS-CMP comprises 2,122 comparison questions created from merged image pairs, where the count difference spans from 1 to 9. Notably, 80% of the comparisons fall within a difference of 1–4. The most common question pairs in OPTICS-CMP involve direct recognition of an object’s taxonomic (56.7%) or physical (22.1%) properties (details in Figure˜10). We also estimated the OPTICS-CNT difficulty for humans by asking two new participants to answer 90 questions about 30 random images from OPTICS-CNT, split across reasoning levels and object properties. Their overall accuracy of 74% indicates that while the task is challenging for humans, they answer most questions correctly.
5 Experimental Setup
Model Selection. To ensure representative evaluation, we include a variety of non-instruction-tuned and instruction-tuned open-source VLMs, as well as a fine-tuned VLM and a closed-source model. As non-instruction-tuned open-source VLMs, we include BLIP-2 (Li et al., 2023) combined with OPT-2.7b and OPT-6.7b (Zhang et al., 2022), and Fuyu-8b (Bavishi et al., 2023). We include the instruction-tuned open-source models BLIP-2 paired with Flan T5-xxl (Chung et al., 2024), Qwen2.5-VL (3B, 7B, and 32B Instruct models) (Bai et al., 2025), and InternVL3 family (8B and 14B variants) (Zhu et al., 2025) prominent on the Hugging Face OpenVLM Leaderboard (Duan et al., 2024), and Gemma 3 (Google et al., 2025b) with 27B parameters. Our study incorporates the SpatialVLM fine-tuned version of Qwen2.5-VL for spatial relations and depth estimation in VQA (Chen et al., 2024), SpaceThinker-Qwen2.5-VL, claimed by the authors to be their most accurate fine-tuned specialized reasoning model. As a closed-source VLM, we include only GPT-4o-mini due to budget considerations (OpenAI et al., 2024). All models are evaluated in a zero-shot setting using standard hyperparameters.
Answer Extraction. For both benchmarks, our VLM prompts facilitate answer extraction by requesting either the total count or a count followed by objects for OPTICS-CNT (see Table˜4), and only “A” or “B” to indicate which value is greater for OPTICS-CMP. To extract answers from OPTICS-CNT (numeric label) and OPTICS-CMP (A or B label), we employ flexible regular expressions for detection and perform exact string matching against ground-truth labels. For OPTICS-CNT, we interpret missing outputs as a prediction of zero, and though during data curation we restrict object counts to 10 per image, we allow model outputs with counts greater than 10 to enable analysis of deviation from ground truth.
Evaluation Metrics. Following established count-based VQA benchmarks (Trott et al., 2018; Acharya et al., 2018), we report multiple evaluation metrics for OPTICS-CNT. Our primary measure is micro accuracy (Acc). We report macro accuracy, which averages the accuracy scores across different counts, ranging from 0 to 10. We also report Root-Mean-Squared Error (RMSE) to quantify the typical deviation between predicted and ground-truth counts. To diagnose model bias, we include mean error, which reveals bias towards over- or under-counting. Furthermore, off-by-N accuracy (cumulative accuracy by tolerance) indicates the proximity of the predictions to the correct count. We also report accuracy per count difference for OPTICS-CMP, expecting an increase in accuracy for larger differences.
6 Results
Model Micro
Acc
($\uparrow$)
Macro
Acc
($\uparrow$)
RMSE
($\downarrow$)
Mean
Error
($\rightarrow 0$)
Off-By-1
($\uparrow$)
Random 9.09 9.09 4.47 0.00 25.62
BLIP-2 OPT (2.7B) 12.87 10.92 3.66 -2.68 35.28
BLIP-2 OPT (6.7B) 16.67 9.68 3.06 -1.56 46.11
Fuyu (8B) 19.91 11.88 2.68 -0.89 49.72
BLIP-2 Flan (11B) 15.93 12.95 3.53 -2.05 42.59
Qwen2.5 (3B) 24.91 23.55 15.30 -1.16 51.85
Qwen2.5 (7B) 38.70 31.57 1.88 -0.62 70.28
Qwen2.5 (32B) 39.91 34.46 1.80 -0.35 72.78
SpaceThinker (3B) 29.63 25.32 2.19 -0.89 62.50
InternVL3 (8B) 37.59 31.30 1.93 -0.70 70.28
InternVL3 (14B) 39.72 29.90 2.08 -0.91 68.61
Gemma3 (27B) 31.76 24.43 2.20 -0.14 65.28
GPT-4o mini 30.37 22.74 15.19 -0.36 64.17
Model average 28.16 22.39 4.62 -1.02 58.28
Human 1 75.56 75.15 4.22 0.64 82.22
Human 2 72.22 67.27 1.63 -0.24 85.56
Human average 73.89 71.21 2.92 0.2 83.89
Table 2: Zero-shot results on OPTICS-CNT. The highest and second-highest model results are highlighted in bold and underlined, respectively. ($\uparrow$)/($\downarrow$) indicates that higher/lower values are better. ($\rightarrow 0$) signifies values closer to zero are better. means the result is attributed to inductive bias.
How well can VLMs accurately count and reason about object properties? The results for OPTICS-CNT in Table 2 show that most models perform above random, but well below human accuracy (28% versus 74% on average). The high RMSE across many of the models indicates substantial variance in predictions, suggesting that precise numerical grounding is challenging, even for strong VLMs. The negative mean error values suggest consistent underestimation and a general bias towards undercounting. We note a severely imbalanced prediction distribution in the outputs of some VLMs (detailed in Figure 14). When allowing for one counting error, the accuracy increases to 58% on average (to 73% for the best model), which indicates that many model errors are near-misses (see Figure 15 for an in-depth analysis of the top-3 best performing VLMs). For all VLMs, the micro accuracy is higher than the macro accuracy because all models exhibit a bias towards lower, more frequent counts, as shown in Figure 5. In contrast, human accuracy is uniform across answer counts, with no apparent frequency bias. We further analyze the sources of errors for both humans and VLMs at the end of this section.
How does counting performance vary across VLMs? Accuracies vary widely across models, ranging from 12.87% to 39.91%. The open-source Qwen2.5-VL-Instruct (32B) is the best-performing model, with the highest exact and off-by-1 accuracies, the lowest RMSE, and the lowest mean error. A small gap separates it from InternVL3 (14B), with the second-highest overall accuracy, and Gemma 3 (27B), which achieves the lowest mean error. The closed-source GPT-4o-mini exhibits moderate performance, with a high RMSE and low mean error. We also note a typical increase in performance with increasing model size within model families, though this may not always hold across architectures. Table 3 shows that Qwen2.5-VL-Instruct (7B) and (32B) demonstrate the strongest performance across nearly all reasoning levels, property dimensions, and image types, while GPT-4o-mini performs consistently worse than most instruction-tuned open-source models.
0112233445566778899101002020404060608080100100Ground Truth CountAccuracy (%)HumanIntern14BQwen7BQwen32B
Figure 5: Accuracy per count (0-10) for human with an average (micro) accuracy of 73.89% and the top-3 best performing models: Qwen2.5-VL-Instruct (32B), InternVL3 (14B), and Qwen2.5-VL-Instruct (7B).
Model Reasoning Complexity Object Properties Image Types Direct Recognition Property Inference Counter- factual P T F R Photo Animated AI Gen. BLIP-2 OPT (2.7B) 10.83 13.06 14.72 11.48 14.41 10.91 13.78 9.17 19.44 10.00 BLIP-2 OPT (6.7B) 16.67 18.06 15.28 16.39 16.52 14.09 19.08 14.17 21.11 14.72 Fuyu (8B) 18.33 22.78 18.61 17.21 17.42 19.55 25.44 16.94 23.89 18.89 BLIP-2 Flan (11B) 18.33 18.61 10.83 14.34 16.82 13.64 18.02 11.67 20.00 16.11 Qwen2.5 (3B) 26.11 29.17 19.44 21.31 29.43 17.73 28.27 20.28 27.78 26.67 Qwen2.5 (7B) 42.78 44.17 29.17 35.66 42.34 32.73 41.70 32.22 40.83 43.06 Qwen2.5 (32B) 39.17 43.06 37.50 33.61 44.44 38.18 41.34 31.94 45.00 42.78 SpaceThinker (3B) 34.44 30.28 24.17 25.00 37.24 19.55 32.51 25.56 31.11 32.22 InternVL3 (8B) 38.89 41.94 31.94 33.20 42.34 26.82 44.17 28.89 43.61 40.28 InternVL3 (14B) 42.78 42.22 34.17 30.33 48.05 36.82 40.28 32.22 48.61 38.33 Gemma3 (27B) 33.89 35.56 25.83 26.23 34.53 28.64 35.69 27.78 36.39 31.11 GPT-4o mini 28.89 33.61 28.61 21.31 34.23 25.00 37.81 24.44 38.89 27.78 Model average 29.34 31.04 24.18 23.83 31.48 23.63 31.50 22.94 33.05 28.49 Human 1 73.33 80.00 73.33 58.33 92.00 71.43 80.00 63.33 83.33 80.00 Human 2 66.67 76.67 73.33 58.33 80.00 76.19 75.00 60.00 83.33 73.33 Human average 70.00 78.33 73.33 58.33 86.00 73.81 77.50 61.66 83.33 76.66
Table 3: Accuracy results for each model on OPTICS-CNT across reasoning levels, object property dimensions (Physical, Taxonomic, Functional, Relational), and image types. The highest and second highest model results are highlighted in bold and underlined, respectively.
How does counting performance vary across reasoning levels? We expect the progression of reasoning complexity recognition << property inference << counterfactual3) to be reflected in corresponding trends in model performance. However, the results among the top-tier VLMs (Qwen2.5-VL-Instruct, InternVL3, Gemma 3, and GPT-4o-mini) reveal a progression property inference << recognition << counterfactual. This suggests that modern VLMs still exhibit poor low-level perceptual capabilities, which aligns with several recent findings: 1) Jiang et al. (2024) showed poor perceptual performance of MLLMs for fine-grained visual analysis tasks, including counting; 2) Zhang et al. (2024a) identified blind spots in MLLMs, including sensitivity to visual quality and image distractors; 3) Yuan et al. (2021) showed perception failures in the MetaVQA framework and highlighted the need for detailed perceptual grounding of VLMs. The lower accuracy on counterfactual questions is indicative of difficulty with changes and hypotheticals, reliance on learned associations, and lack of reliable grounding Li et al. (2024). Humans also perform best on inference questions, but struggle less than VLMs with counterfactual reasoning.
How does counting accuracy vary across object properties? Among the object property dimensions, most models (as well as humans) have a stronger grasp of taxonomic and relational questions than of physical and functional ones. This finding is in line with recent work that points out the struggle of LLMs with physical object attributes (Wang et al., 2023), physical interactions (Aroca-Ouellette et al., 2021), functional requirements (Qasemi et al., 2022), and atypical affordances (Tian et al., 2025). We note an improvement for the fine-tuned SpaceThinker-Qwen2.5-VL (3B) model compared to its non-fine-tuned counterpart, Qwen2.5-VL (3B), across the four object property dimensions. This shows that, in addition to relational questions, other object properties also benefit from spatial tuning of Qwen2.5 (3B). We also observe that the InternVL3 models have a better grasp of taxonomic and relational questions compared to other models. The Qwen2.5-VL (7B) and (32B) variants demonstrate consistent performance across all properties, with particular strength in physical and functional questions. In contrast, the non-instruction-tuned models struggle across all properties, underscoring their weakness in visual object reasoning.
How does counting performance vary across image types? We observe a consistently higher accuracy for AI-generated and animated images, compared to photographic images. We hypothesize that this is due to the visually cleaner, less cluttered nature of AI-generated and animated images compared to the complex, noisy real-world scenes in photographic images. This finding aligns with previous studies reporting similar gaps between synthetic and real-world data Johnson et al. (2017); Hudson and Manning (2019). Moreover, we note that animated images are easier for most models than AI-generated images, possibly because the latter may have poor focus and include implausible objects, such as a half-cut glass floating in the air. The trend between the image difficulty animated << AI-generated << photographic is also clear in humans, which indicates that questions about photographic images may be the hardest and possibly contain remaining ambiguity.
How well do VLMs answer comparison questions, and is this sensitive to the count differences? Based on results from OPTICS-CNT, we evaluate only the instruction-tuned models on OPTICS-CMP. This experiment confirms that VLMs struggle with counting objects, even in this relative setting. Namely, VLMs’ accuracy remains only slightly higher than random, at 58% (details in Table 9). Qwen2.5-VL (32B) remains the overall best-performing model, followed by InternVL3 (8B), which outperforms its larger (14B) variant. This pattern does not hold within the Qwen2.5-VL family, where we instead observe a steady increase in performance as model size increases. Our hypothesis that VLMs would show higher gains on larger count differences is examined in Figure 6. While some models clearly follow this trend, we note that the model performance across the count differences remains broadly stable. These findings confirm those from the absolute-counting task but also demonstrate that the failure modes of humans and VLMs may differ. We next dive deeper into the causes for errors in both humans and VLMs.
123456-940405050606070708080Ground Truth Count DifferenceAccuracy (%)Gemma27BGpt4ominiIntern14BIntern8BQwen32BQwen3BQwen7B
Figure 6: Accuracy per count difference (1-9) for instruction-tuned VLMs on OPTICS-CMP. Counts (6-9) are grouped as one.
Figure 7: Examples of errors made by humans and the model. Left: Both the model and the human made a perception error. Right: The human made a semantic ambiguity error, while the model made an unnatural error.
How similarly do humans and VLMs make mistakes? To investigate this, we selected 25 questions for which both the best VLM (Qwen2.5-VL-32b) and at least one human judge gave an incorrect answer relative to the ground truth. We defined four categories of errors: 1) Semantic ambiguity errors caused by unclear semantic constraints in the question, e.g., should an object that is half wood and half metal be counted as “metal”? 2) Visual ambiguity errors caused by ambiguous visual information in the image, e.g., should we count a chair if only part of its legs is visible and not the whole object? 3) Perception errors - honest mistakes due to overlooking or misperceiving information, e.g., not noticing a small object or miscounting the number of items. 4) Unnatural errors that are highly unlikely for a human to make, e.g., counting a glass as a metal object or counting an object that clearly does not exist.
We assigned each incorrect answer from the model and the human judge to one or more of these four categories. We observe the following (full results in Table 10): a) Humans do not make unnatural mistakes (0/25), whereas the model makes many of them (9/25). This is the most significant difference between humans and the model, which may stem from hallucination/illusion-driven errors in models, as evidenced by diagnostic suites (Guan et al., 2024). An example is shown in Figure 7 (right), where the human counted all objects that could be considered metal, while the model counted items like the glass and the cutting board, which are clearly not metal. b) For humans, the majority of mistakes (16/25) come from semantic ambiguity, not knowing whether something should be counted or how to interpret a specific condition. This finding is consistent with our analysis of image types, which concluded that photographic images may be most challenging due to remaining ambiguity in the questions or in the visual content. c) The second most common source of human error is perception errors (9/25), often due to not examining the image carefully enough and missing objects. Perception errors are similarly frequent in humans and the model (9 and 8, respectively). An example is shown in Figure 7 (left), where both the human and the model made the same mistake: they missed the turtle inside the water.
7 Conclusions
This paper developed a commonsense evaluation framework for evaluating object property reasoning in images. The framework considered three distinct image types (photographic, animated, and AI-generated), three reasoning levels (direct recognition, property inference, counterfactual), and four property types (physical, taxonomic, functional, relational). We instantiated this framework on two VQA benchmarks: OPTICS-CNT, with over 1,000 counting questions across 360 images, and OPTICS-CMP, with over 2,000 comparison questions across as many images. Experiments with 12 state-of-the-art VLMs revealed low counting accuracy, with most models performing far below human performance and the best achieving only 40% accuracy. VLMs struggle particularly with realistic (photographic) images, counterfactual reasoning about physical and functional properties, and infrequent counts over five. Humans also struggle with photographic images and physical properties but excel at counterfactual questions and have no frequency effects on counts. Interestingly, when asked to compare images based on their counts, most VLMs performed better for higher differences, but their performance on very large differences remained relatively modest. Analyzing closely the errors made by humans and VLMs revealed that, while both make a sizeable proportion of perceptual mistakes, VLMs are dominated by unnatural errors, whereas humans struggle with semantic ambiguity. This indicates that VLMs experience limitations on a combination of perception and reasoning abilities and unexpected hallucinations.
Future work should address three emerging limitations. First, generative AI still suffers from limited accuracy and diversity in autonomous question generation, causing scaling up object reasoning benchmarks to require substantial human effort. Our attempt to generate a scalable version based on richly annotated data (§B.5) showed that automatically creating such benchmarks remains challenging. Novel approaches, such as metamorphic testing Yuan et al. (2021), are necessary to grow OPTICS-CNT systematically in size and coverage of additional object properties, question types beyond counting and comparison, and open-ended answers. Second, since annotators are free to design their own questions, they may introduce semantic ambiguity, leading to disagreements (e.g., whether zebras are considered black animals). Although we addressed ambiguity through an explicit quality assurance step that analyzes sources of disagreement and translates them into more precise guidelines, some disagreements remain. Future work should develop generalizable and well-understood guidelines for object abstraction and reasoning. Finally, while OPTICS-CNT and OPTICS-CMP point to fundamental limitations of popular and diverse VLMs, the set of baselines is inherently incomplete. Future work should explore additional specialized reasoners beyond SpatialVLM, e.g., a referring expression counting method to address compositional attribute filters and counting Dai et al. (2024), mixtures of experts (Lin et al., 2024), derivation of object schemas Hsu et al. (2024), and methods for populating situational scene graphs following predefined slots Sugandhika et al. (2025).
Appendix A Appendix: Technical Details
Category Prompt Template Non-Instruction-tuned models {IMG}
Question: {question} Provide only the total number. Answer:
Instruction-tuned models {IMG}
{question} Your response MUST be in the following format and nothing else: <<NUMBER>> [<<OBJECT1>>, <<OBJECT2>>, <<OBJECT3>>, …]
Table 4: Prompt template used to extract output in a specific format from the model for non-instruction-tuned and instruction-tuned models. {IMG} and {question} are the placeholders for the image and question, respectively.
Image Domain Prompt Example Campsite settings (Outdoors) Create image Generate a photorealistic dusk campsite scene with 5 tents, a few of them illuminated from within, a controlled campfire with 3 people sitting on logs around it and a dog sleeping nearby, hiking backpacks leaning against a nearby tree, a lantern hanging from a branch, and stars beginning to appear in the darkening sky. Include a mountain silhouette in the background. Strictly keep the count of total objects in the image to a maximum of 10. Gym (Indoors/Household) Create image Generate a photorealistic image of a gym scene with sports equipment. Let the gym contain 3 treadmills, 2 elliptical machines, a few kettlebells on the floor, 2 resistance bands hanging on the wall, and a bench press rack. Let it be a realistic setting with 2 humans working out, one person on the treadmill and another person on the elliptical machine. The gym can have a clock and TV hanging on the wall somewhere showing a workout video. Include large windows with natural light and a small plant in the corner. Strictly keep count of total objects in the image to a maximum of 10. Beach cleanup (Outdoors) Generate a photorealistic image of a beach cleanup scene with 4 volunteers in bright t-shirts collecting trash at sunset in the backdrop. Let the trash include these objects: 5 glass bottles, 6 trash bags (one in each hand of a volunteer and 2 placed near the ground), a collection bin, reusable gloves, and garbage pickers. Let it be a realistic setting, having the ocean with gentle waves in the background. Strictly keep the count of total objects in the image to a maximum of 10. Garage (Neutral) Create image Create a photorealistic image of a vehicle garage scene during a vehicle maintenance session. Show 3 cars, one car on a hydraulic lift with its hood open, 2-3 essential tools spread on a workbench, an oil change in progress, spare parts in organised containers, and a service manual. Let the setting be realistic with ceiling lights for the garage, let the car being serviced be in good condition, and let the other two cars in the background be scrap and old, which need to be rebuilt with loose tires leaning next to them. Let one person work on the car that is being serviced and the other person in the backdrop next to the other 2 cars. Strictly keep count of total objects in the image to a maximum of 10. Kitchen (Indoors/Household) Create image Generate an image of a kitchen countertop during the preparation of a complex dish. Show fresh ingredients arranged in 3 small ceramic bowls, a wooden cutting board with chopped herbs, a professional knife set, a simmering pot on the stove, a recipe book, a glass filled with water, and 2 people collaborating on the cooking. Include details like steam, a kitchen sink next to the stove, and ambient kitchen lighting. Let it be a realistic setting. Strictly keep the count of total objects in the image to a maximum of 10.
Table 5: Prompt examples used to generate AI images across different domains, along with corresponding tags (shown in brackets) used in constructing the MacGyver dataset. Prompts with the Create Image tool are generated using GPT-4o, while those without it are created with Grok 3.
Indoors/Household Neutral Outdoors bedroom tools beach cleanup home setup tech camping kitchen garage construction library classroom gardening laboratory biking meeting room market gym zoo salon urban wardrobe park picnic farm driveway bustop
Table 6: The 15 domains (i.e., locations and activities) taken from the MacGyver dataset (Tian et al., 2025), extended with 11 of our own, for a total of 26 unique image domains. They are broadly divided into Indoors/Household, Neutral, and Outdoors.
A.1 Prompt Strategies
A.1.1 Question Generation
We pass a standard prompt template to any of the three MLLMs: GPT-4o, Claude 3.7 Sonnet, and Gemini 2.0 Flash to generate questions, which are then used to inspire humans to generate new questions manually. The template is as follows:
{IMG} Generate three questions per prompt for this image with answers for each prompt. Prompt 1 deals with physical/taxonomic properties; prompt 2 deals with functional/relational properties; and prompt 3 deals with counterfactual reasoning. All need to be counting questions; be creative. You do not need to stick to the same questions.
A.1.2 Output Formatting
Table 4 displays the prompt templates used on instruction-tuned and non-instruction-tuned models to extract outputs in a specific format. Instruction-based models are also asked for a list of objects detected and considered for the count, which is currently not directly evaluated.
A.1.3 AI-Generated images
Table 5 shows the prompt examples used while generating AI-based images for different image domains.
A.2 Image Domains
During image collection (Figure 2), to direct our search, we made use of image domains inspired by the tags used in curating the MacGyver dataset Tian et al. (2025). The three broad categories in this dataset are: indoors/household (e.g., kitchen), neutral (e.g., garage), and outdoors (e.g., zoo). Table 6 shows the list of 26 unique image domains used for the curation of OPTICS. The process of selecting a domain was influenced by the chance of a particular domain having a greater expected variety of objects. The domain tags used for the MacGyver dataset were plentiful, out of which 15 were in common, namely: bedroom, kitchen, library, meeting room, gym, wardrobe, garage, classroom, beach cleanup, camping, construction, gardening, zoo, park, and farm. The remaining 11 tags resulted from our exploration.
A.3 Benchmark Examples
A.3.1 Image Types
Figure 8 shows additional examples from OPTICS-CNT for three image types: Photographic, Animated, and AI-generated.
Figure 8: Examples from OPTICS-CNT, for three image types: Photographic (top), Animated (middle), and AI-generated (bottom). Text highlighted in blue is information stored in the benchmark and not passed to the model during evaluation. GT stands for ground truth, RC for reasoning complexity, and OP for object properties.
A.3.2 Object Property Dimensions
Table 7 shows examples of questions present per object property dimension.
Dimensions Question Examples Physical How many objects made of wood are present? How many objects are present that are transparent? How many objects in the background are present that have legs? Taxonomic How many mammals are visible in the image? How many furniture items are present in the room? How many tools are visible in the image? Functional How many objects with the primary purpose of illumination can be seen? Count the number of breakable items? Count the number of items that are battery powered? Relational How many objects are visible that are attached to the wall or ceiling? How many reptilian couples, at maximum, are present?
Table 7: Question examples for the different object property dimensions: physical, taxonomic, functional, and relational.
A.3.3 Reasoning Complexity
Table 8 shows examples of the questions asked at the different reasoning levels. Direct recognition has questions from the physical and taxonomical object property dimensions; property inference has questions from the functional and relational dimensions; and counterfactual reasoning has questions from any of the four property dimensions.
Complexity Dimension Question Example Direct Recognition physical How many objects made of wood are present? taxonomic How many mammals are visible in the image? Property Inference functional Count the number of breakable items? relational How many objects are visible that are attached to the wall or ceiling? Counterfactual physical If one of the metal objects were replaced by a wooden object, how many wooden objects would be there in the image? taxonomic If one person leaves the cleaning group, how many mammals would remain? functional If the two bedside lamps were removed, how many objects are present that need electricity? relational If the signages were removed, how many objects would be present that hang from the ceiling?
Table 8: Question examples for the different reasoning complexity levels that is direct recognition, property inference, and counterfactual.
A.3.4 OPTICS-CMP
Figure 9 shows an example from OPTICS-CMP where we merge two images from OPTICS-CNT into a single image. Figure 10 visualizes the sample distribution across count-difference levels and question categories.
Figure 9: Example from OPTICS-CMP with a count difference of 2.
Figure 10: Sample Distribution of OPTICS-CMP.
A.4 Ground Truth Distribution
For the total number of 1,080 questions, the ground truth distribution for all the counts ranging from 0 to 10 is shown in Figure 11.
Figure 11: Ground Truth distribution for all counts ranging from 0 to 10.
A.5 Examples of Model Errors
OPTICS requires models to have good low-level perception and the ability to perform fine-grained analysis. A case study on the errors models typically make is visually depicted using an image of photographic type. Figure 12 shows an instance from OPTICS-CNT, with the ground truth values provided after the question in the yellow cell and the depiction of object counts through green bounding boxes. The bounding boxes are manually annotated on the images using Canva. [^2] https://www.canva.com/ To showcase the output of the model, we consider the open-source model, Gemma 3 (27B), for this study. Figure 13 displays the model’s predicted values and the probable object selection, with green (correct) and red (incorrect) bounding boxes based on detailed reasoning explicitly requested for this image.
For the first question, the model outputs 5, as it misses the sparrow that is partially occluded but still visible to the human eye. In the second question, the model outputs 6, as it tends to over-count the row of chairs due to the challenging angle at which the picture is captured. In the final question, the model outputs 2, reasoning that there are three birds (excluding the partially visible one) on the right sitting together on one railing, leaving two birds on the left. But in the ground truth, the bird positioned on the extreme right is sitting on a different metal frame. Based on the model’s answer to the first question, its reasoning is consistent with the third question’s answer.
The model fails to detect objects that are partially hidden, indicating weak low-level perception. It misinterprets perspective and geometric layout and struggles to distinguish objects that are closely clustered together. Despite the incorrect answers, the model’s internal reasoning is logically consistent. This study highlights the importance of evaluating not only the final count, but also the quality of visual grounding and traceability of reasoning for tasks like those in OPTICS, providing potential for future work.
Figure 12: Questions follow the reasoning complexity order, direct recognition to counterfactual (left to right) for a photographic image with the ground truth values and the object counts visually depicted by green bounding boxes.
Figure 13: Gemma 3 model output values and the visual depiction of possible correct and incorrect detections.
Appendix B Appendix: Complementary Results
Figure 14: BLIP2 model series inductive bias plot, consistently defaulting to specific output values.
Figure 15: Cumulative accuracies for tolerance levels 0, 1, and 2. The top-performing models from the following categories are shown: open-source non-instruction VLM (Fuyu-8B), open-source instruction VLM (Qwen2.5-VL-32B-Instruct), closed-source VLM (GPT-4o mini), the human evaluators average and the random chance baselines for tolerance levels 0 (9.09%), 1 (25.62%), and 2 (40.50%).
B.1 Inductive Bias
The BLIP2 model series consistently defaults to specific output values, suggesting the presence of inductive bias in the models. The OPT 2.7B variant goes for the count 1, the OPT 6.7B variant produces either 1 or 3 as the output, and the Flan T5 xxL gives an output of 0 or no value or an answer not abiding by the required format (numeric value), resulting in a final output of 0 set according to our answer extraction method. The frequency distribution of these outputs is shown in Figure 14. We tried adjusting prompts and decoding parameters, such as the number of maximum new tokens, but none helped mitigate the issue, highlighting the potential inductive biases present in models (Wang and Wu, 2023; Kraaijveld et al., 2025).
B.2 Cumulative Accuracies with Tolerance
The soft off-by-1 accuracy metric, which has a baseline of 25.62%, reveals relatively better improvements (47.16% gap with the best model) among the larger instruction-tuned models. Even though the models predictions are often nearly correct, this metric highlights the difficulty in precise numeric predictions by VLMs. The cumulative accuracy for tolerance values 0, 1, and 2 is shown in Figure 15 for the top-performing models in their categories and the average of the human evaluators.
B.3 OPTICS-CMP results
Table 9 shows the results for seven instruction-tuned models evaluated on OPTICS-CMP.
Model Micro
Acc
($\uparrow$)
Macro
Acc
($\uparrow$)
Qwen2.5 (3B) 55.14 56.30
Qwen2.5 (7B) 57.54 58.10
Qwen2.5 (32B) 68.33 72.39
InternVL3 (8B) 62.02 64.35
InternVL3 (14B) 46.84 64.19
Gemma3 (27B) 54.52 64.48
GPT-4o mini 60.34 64.45
Model average 57.81 63.46
Table 9: Zero-shot results on OPTICS-CMP. The micro accuracy and macro accuracy (average across count differences (1-9)) are reported. The highest and second-highest model results are highlighted in bold and underlined, respectively. ($\uparrow$) indicates that higher values are better.
B.4 Error Analysis of Humans and Qwen
The raw results of this analysis are provided in Table 10.
Semantic
ambiguity
Visual
ambiguity
Perception
errors
Unnatural
errors
Human 16 3 9 0
Qwen2.5 (32B) 8 1 8 9
Table 10: Results from analyzing the errors made by humans and the model.
Figure 16: Left: A successful example of a question generated using Visual Genome. Right: An unsuccessful example, where multiple objects are assigned to a single tree and a single patch of grass, leading to an incorrect answer.
B.5 Visual Genome-based Dataset Curation
Figure 16 shows two examples of question-answer pairs generated automatically from Visual Genome. We pose the following question: Can densely annotated datasets be used to scale up OPTICS-CNT automatically? Visual Genome (VG) Krishna et al. (2017) is a popular dataset with high-quality structured connections between image regions, language descriptions, and WordNet Miller (1995) synsets. To investigate whether the annotations of VG can support the generation of large-scale instantiations of OPTICS, we devised a procedure that generates taxonomic questions by manipulating the WordNet synsets. This experiment exposed three limitations: lack of contextualization of classes (e.g., kitchen utensil instead of a bowl in Figure 16-left), incorrect questions or answers (e.g., about uncountable objects like grass in Figure 16-right), and classes in the WordNet hierarchy that are not overly specific (e.g., derby horse race) or generic (e.g., entity). Thus, while every image in VG enables many taxonomic questions to be created, it remains challenging to automatically distinguish high- from low-quality question-answer pairs. More fundamentally, creating question-answer pairs for other object properties (e.g., functional) and reasoning types (e.g., counterfactual) is hindered by the absence of annotation. While LLMs can be employed to generate questions for these categories, we established that these questions and answers are often incorrect or ambiguous (e.g., suggesting that buildings, doors, and windows can all provide shelter). In conclusion, we find that densely annotated image datasets remain insufficient to generate high-quality large-scale versions of OPTICS. Yet, they could potentially be used to speed up human annotation by providing an initial set of samples or for distant supervision of VQA models.